A balloon originally had a volume of 185 mL at 102C. What is the new temperature of the gas in the balloon if the volume of the balloon changes to 347 mL?
Answer #1

Charles law volume of a gas the absolute At constant pressure, the is directly proportional to temperature va T V=KT k is calV = 185 ml = 102+273= 375k V₂ = 347md T,=? 185ml 375 K 347ml 375k X 347 ml I= 185 ml | Tg = 703-3783 78378 k